# Progressive-Resolution Policy Distillation: Leveraging Coarse-Resolution Simulations for Time-Efficient Fine-Resolution Policy Learning

### v3

@@ Abstract
In earthwork and construction, excavators often encounter large rocks mixed with various soil conditions, requiring skilled operators. This paper presents a framework for achieving autonomous excavation using reinforcement learning (RL) through a rock excavation simulator. In the simulation, resolution can be defined by the particle size/number in the whole soil space. Fine-resolution simulations closely mimic real-world behavior but demand significant calculation time and challenging sample collection, while coarse-resolution simulations enable faster sample collection but deviate from real-world behavior. To combine the advantages of both resolutions, we explore using policies developed in coarse-resolution simulations for pre-training in fine-resolution simulations. To this end, we propose a novel policy learning framework called Progressive-Resolution Policy Distillation (PRPD), which progressively transfers policies through some middle-resolution simulations with conservative policy transfer to avoid domain gaps that could lead to policy transfer failure. Validation in a rock excavation simulator and nine real-world rock environments demonstrated that PRPD reduced sampling time to less than 1/7 while maintaining task success rates comparable to those achieved through policy learning in a fine-resolution simulation.
@@
Index Terms:
Simulation Resolution, Excavation, Policy Distillation, Reinforcement Learning
@@ I Introduction
[…]
@@ III-A Reinforcement Learning
where $A(s,a)=Q(s,a)-V(s)$ is the advantage function. The value function $Q(s,a;\psi)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\big{|}s_{0}=s,a_{0}=a\right]$ represents the expected cumulative reward when action $a$ is taken in state $s$. The value function $V(s;\phi)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\big{|}s_{0}=s\right]$ represents the expected cumulative reward starting from state $s$ and following policy $\pi(s,a;\theta)$. These functions are typically approximated using function approximators such as neural networks. Here, $\theta$, $\phi$, and $\psi$ denote the parameters of the policy $\pi$, the value functions $V,Q$, respectively.
@@ IV-A2 Scheduling Setups
Algorithm 1 Progressive-Resolution Policy Distillation with PPO # Set parameters described in Table I
@@
for $n=1,2,...,N$ do
       # Initialize iteration num $i\!=\!0$, parameter update num $k\!=\!0$
@@ V-D Variable Resolution Simulator
[…]
Error rate of rock in bucket []delimited-[][$\cdot$][ ⋅ ] 0 0.2 0.2
[…]
Bucket torque weight (XYZ) []delimited-[][$\cdot$][ ⋅ ] 1 0.8 1.2
[…]
Friction coefficient []delimited-[][$\cdot$][ ⋅ ] 1 0.8 1.2
[…]
Total soil volume $[\textrm{mm}^{3}]$ $125^{3}$ $120^{3}$ $130^{3}$
Total rock volume $[\textrm{mm}^{3}]$ $50^{3}$ $45^{3}$ $55^{3}$
@@
Trained Resolution $\Delta$ [mm] 70 60 50 40 30 20 10
Trained Resolution [%] $99.4\pm 2.6$ $99.1\pm 3.1$ $98.9\pm 2.3$ $98.3\pm 1.9$ $97.5\pm 5.6$ $96.9\pm 7.2$ $95.8\pm 6.4$
@@ VI-A Common Settings
Figure 6: Summaries of analysis: (Left) Performance comparison of PRPD components for (a) different patterns of policy mixture rate, (b) different numbers of grid interval, and (c) different numbers of target middle success rate. The box plots represent the sample numbers at which the policies achieve the target success rate $\hat{\tau}$ for each resolution. The dashed lines denote the total learning time from the initial resolution to the plotted resolution. $\alpha$ denotes the policy mixture rate, which is dynamically updated only in $\alpha=\text{opt.}$.. (Right) (d) different scaling parameters of loss function in policy learning required to reach $\hat{\tau}$. Each curve and point of (a), (b), (c), and (d) plots the mean and variance per sample over five experiments (each box plot plots the mean). The learning sample is the value until the target success rate $\hat{\tau}$ is reached at the corresponding resolution (for Ours in (d), it is the summation value until $\hat{\tau}$ is reached at the final resolution $\Delta=10$). * mean $p<0.05$.
@@ VI-B2 Results
[…]
@@ VI-C2 Results
As shown in Fig. 6, $\alpha=\text{opt.}$. achieves the shortest learning time compared to fixed mixture rates. Additionally, $\alpha=1$, where previous policies $\pi^{(n-1)}$ are used without optimization, outperforms $\alpha=0$, which does not use previous policies. These results confirm that the conservative policy transfer scheme enables learning with fewer samples.
@@
TABLE V: Sim-to-real experiment of rock excavation: This experiment evaluated nine types of rocks (five differently shaped wooden blocks and three different-sized 3D-printed artificial rocks as shown in rock images) on three types of ground (for each element, three numbers correspond to each environment, silica sand, kinetic sand, and plastic pellets in order). “Scale” indicates approximate rock size. The first-row block shows “Ours” (PRPD) and “w/o DR” (PRPD without DR). The second-row block shows PPO at a fixed resolution. Each value indicates the success number per 20 trials, with the last column showing the average rate. The 20 trials test five policies at four positions (0, 60, 120, and 180 degrees) on a 100-mm radius circumference relative to the bucket’s initial position. Task success is defined as rocks being held in the bucket and the bucket being off the ground.
@@ VI-G Sim-to-Real Policy Transfer
[…]

### v4

@@ Abstract
In earthwork and construction, excavators often encounter large rocks mixed with various soil conditions, requiring skilled operators. This paper presents a framework for achieving autonomous excavation using reinforcement learning (RL) through a rock excavation simulator. In the simulation, resolution can be defined by the particle size/number in the whole soil space. Fine-resolution simulations closely mimic real-world behavior but demand significant calculation time and challenging sample collection, while coarse-resolution simulations enable faster sample collection but deviate from real-world behavior. To combine the advantages of both resolutions, we explore using policies developed in coarse-resolution simulations for pre-training in fine-resolution simulations. To this end, we propose a novel policy learning framework called Progressive-Resolution Policy Distillation (PRPD), which progressively transfers policies through some middle-resolution simulations with conservative policy transfer to avoid domain gaps that could lead to policy transfer failure. Validation in a rock excavation simulator and nine real-world rock environments demonstrated that PRPD reduced sampling time to less than 1/7 while maintaining task success rates comparable to those achieved through policy learning in a fine-resolution simulation. Additional videos and supplementary results are available on our project page: https://yuki-kadokawa.github.io/prpd/
@@ I Introduction
We propose a novel policy learning framework called Progressive-Resolution Policy Distillation (PRPD) for achieving time-efficient policy learning of fine-resolution simulations by utilizing coarse-resolution simulations effectively. PRPD progressively improves simulation resolution while learning and transferring policies at each stage. PRPD uses a conservative policy transfer scheme to regularize policies, stabilizing policy transfer across different resolution simulations having different behaviors. We demonstrated the effectiveness of PRPD by constructing a variable-resolution rock excavation simulator using Isaac Gym [21]. This paper demonstrated a 7-fold improvement in learning time efficiency (PRPD: 90 minutes, policy learning in the finest-resolution simulation: 600 minutes), resulting in an 8-hour time gap. This difference is likely to become more significant with increasing state-action space complexity or task difficulty. Furthermore, in evaluations of the learned policies within a real-world environment containing nine types of rocks, PRPD achieved rock excavation in approximately $90\text{\,}\%$ of scenarios.
@@ III-A Reinforcement Learning
where A(s,a)=Q(s,a)V(s)A(s,a)=Q(s,a)-V(s) is the advantage function. The value function $Q(s,a;\psi)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\big|s_{0}=s,a_{0}=a\right]$ represents the expected cumulative reward when action $a$ is taken in state $s$. The value function $V(s;\phi)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\big|s_{0}=s\right]$ represents the expected cumulative reward starting from state $s$ and following policy $\pi(s,a;\theta)$. These functions are typically approximated using function approximators such as neural networks. Here, $\theta$, $\phi$, and $\psi$ denote the parameters of the policy $\pi$, the value functions $V,Q$, respectively.
@@ IV-A2 Scheduling Setups
Algorithm 1 Progressive-Resolution Policy Distillation with PPO
# Set parameters described in Table I
@@
for n=1,2,,Nn=1,2,...,N do
  # Initialize iteration num i=0i\!=\!0, parameter update num k=0k\!=\!0
@@ V-D Variable Resolution Simulator
Obs. of Rock pos. noise (XY) $\textrm{mm}]$ 0 -25 25
Obs. of Rock pos. bias (XY) $\textrm{mm}]$ 0 -25 25
Error rate of rock in bucket $\cdot$] 0 0.2 0.2
Ground height bias $\textrm{mm}]$ 0 -25 25
Init. bucket pos. bias (XYZ) $\textrm{mm}]$ 0 -300 300
Init. rock pos. bias (XYZ) $\textrm{mm}]$ 0 -30 30
Bucket torque weight (XYZ) $\cdot$] 1 0.8 1.2
External force to rock $\textrm{N}]$ 0 0 1
Friction coefficient $\cdot$] 1 0.8 1.2
Total soil mass $\textrm{kg}]$ 3 2.7 3.3
Total rock mass $\textrm{kg}]$ 1 0.8 1.2
Total soil volume $\textrm{mm}^{3}]$ $125^{3}$ $120^{3}$ $130^{3}$
Total rock volume $\textrm{mm}^{3}]$ $50^{3}$ $45^{3}$ $55^{3}$
@@
Trained Resolution $\Delta$ [$m$] 70 60 50 40 30 20 10
Trained Resolution [%] $9.4\pm 2.6 99$$9.1\pm 3.1$ $98.9\pm 2.3$ $98.3\pm 1.9$ $97.5\pm 5.6$ $96.9\pm 7.2$ $95.8\pm 6.4$
@@ VI-A Common Settings
Figure 6: Summaries of analysis: (Left) Performance comparison of PRPD components for (a) different patterns of policy mixture rate, (b) different numbers of grid interval, and (c) different numbers of target middle success rate. The box plots represent the sample numbers at which the policies achieve the target success rate $\hat{\tau}$ for each resolution. The dashed lines denote the total learning time from the initial resolution to the plotted resolution. $\alpha$ denotes the policy mixture rate, which is dynamically updated only in $\alpha=\text{opt.}$. (Right) (d) different scaling parameters of loss function in policy learning required to reach $\hat{\tau}$. Each curve and point of (a), (b), (c), and (d) plots the mean and variance per sample over five experiments (each box plot plots the mean). The learning sample is the value until the target success rate $\hat{\tau}$ is reached at the corresponding resolution (for Ours in (d), it is the summation value until $\hat{\tau}$ is reached at the final resolution $\Delta=10$). * mean $p<0.05$.
@@ VI-B2 Results
As shown in Fig. 5, PRPD learns control policies in less than one-seventh of the time needed to reach the highest success rate compared to fixed resolution learning. Coarser resolutions result in faster convergence, with a 20-fold difference in learning time between $\Delta=70$ and $\Delta=10$, but they also reduce performance by about $35\text{\,}\%$. Also, Table IV shows that each policy achieves over $95\text{\,}\%$ success rate when evaluated at its own training resolution, while the success rate at $\Delta=10$ gradually decreases as the training resolution becomes coarser. PRPD converges in less than one-fourth of the time compared to simultaneous learning across all resolutions. These results demonstrate that the proposed resolution scheduling framework achieves the fastest policy learning.
@@ VI-C2 Results
As shown in Fig. 6, $\alpha=\text{opt.}$ achieves the shortest learning time compared to fixed mixture rates. Additionally, $\alpha=1$, where previous policies $\pi^{(n-1)}$ are used without optimization, outperforms $\alpha=0$, which does not use previous policies. These results confirm that the conservative policy transfer scheme enables learning with fewer samples.
@@
TABLE V: Sim-to-real experiment of rock excavation: This experiment evaluated nine types of rocks (five differently shaped wooden blocks and three different-sized 3D-printed artificial rocks as shown in rock images) on three types of ground (for each element, three numbers correspond to each environment, silica sand, kinetic sand, and plastic pellets in order). “Scale” indicates approximate rock size. The first-row block shows “Ours” (PRPD) and “w/o DR” (PRPD without DR). The second-row block shows PPO at a fixed resolution. Each value indicates the success number per 20 trials, with the last column showing the average rate. The 20 trials test five policies at four positions (0, 60, 120, and 180 degrees) on a 100-$m$ radius circumference relative to the bucket’s initial position. Task success is defined as rocks being held in the bucket and the bucket being off the ground.
@@ VI-G Sim-to-Real Policy Transfer
This section validates the sim-to-real policy transfer across diverse real-world rock-excavation environments, including silica sand (low viscosity), kinetic sand (high viscosity), and plastic pellets (blocky shape and large grain size). Table V shows the evaluation of the transfer policies. The experiment shows that the success rate increases as the resolution improves in the real-world environment for fixed resolution policies, as was demonstrated with the simulation results. PRPD achieves performance comparable to fine-resolution fixed policy learning, indicating no adverse effects from using coarse-resolution simulations and successful sim-to-real transfer. We also evaluated the differences in policy performance during sim-to-real transfer depending on the presence of DR parameters. Table V shows that policy performance dropped to approximately $50\text{\,}\%$ without DR. This indicates a reality gap between the simulation and real-world environments, mitigated by applying DR. These results demonstrate that the variable-resolution rock excavation simulator and the proposed policy learning framework achieve approximately $90\text{\,}\%$ success rates for nine types of real rock environments.